OpenStudy (anonymous):

What is 5/18 + (-1/6) ? A. 1/9 B. 2/3 C. 4/9 D. 1/3 Help Please !!!!!!!!!!!!

OpenStudy (anonymous):

5/18=0.27777777777

OpenStudy (anonymous):

(-1/6)=-0.16666666666

OpenStudy (anonymous):

so 0.27777777777+-0.16666666666=0.1111111

OpenStudy (anonymous):

which is 1/9

OpenStudy (anonymous):

Find a common denominator. 6*3=18 1*3=3 So, \[\frac{ 5 }{ 18 }-\frac{ 3 }{ 18 }=?\]

OpenStudy (anonymous):

OrangeMaster that would be 2/18 which is 1/9

OpenStudy (anonymous):

do you get it shernandez03?

OpenStudy (anonymous):

Yes, but it is more simple that using long decimals.

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

_ though really it is actually supposed to be infinitley ex0.27

OpenStudy (anonymous):

ya thx guys

OpenStudy (anonymous):

your welcome